OpenStudy (anonymous):

What is the slope of the line that passes through the points (-4, 6) and (2, 4)?

OpenStudy (freckles):

\[slope=m=\frac{\Delta y}{\Delta x}=\frac{y_2-y_1}{x_2-x_1}\]

OpenStudy (anonymous):

\[\large slope = \frac{y_2 - y_1 }{x_2 - x_1}\]

OpenStudy (freckles):

Use this substitution : \[(x_1=-4,y_1=6) \text{ and } (x_2=2,y_2=4)\]

OpenStudy (anonymous):

thus, 4-6/2-(-4) = -2/6 = -1/3

OpenStudy (anonymous):